Twin brothers, Andy and Brian, can mow their grandparent's lawn together in 72 minutes. Billy could mow the lawn by himself in 17 more minutes than it would take Andy. How long would it take each person to mow the lawn alone?
Found 2 solutions by jorel1380, addingup:
Answer by jorel1380(3719)

Let a be for Andy, and b be for Brian. Then:
1/a + 1/b=1/72
Brian takes 17 more minutes than Andy, working alone so b=a+17. Then:
1/a + 1/a+17=1/72
72/a + 72/a+17=1
72(a+17)+72a=aČ+17a
aČ-127a-1224=0
(a+9)(a-136)=0
a=-9 or 136
Using the positive value, we get Andy's time working alone as 136 minutes, and Brians' time as 153 minutes. ☺☺☺☺

Answer by addingup(3677)

So you have three people mowing, the twins Andy and Brian, and Billy. All you can say is that Billy is slower than Andy by 17 minutes. There is information missing to be able to answer your question